OpenStudy (anonymous):

MEDAL & FAN The table below represents the distance a person runs as a function of time. The average rate of change of the function between x = 10 to x = 16 is __________ m/s and represents the person's speed. Answer for Blank 1: __________

OpenStudy (anonymous):

Time Distance (seconds) (meters) x f(x) 10 60 13 78 16 96 19 114

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} x&y\\ \hline\\ 10 & 60 \\ 13& 78\\ 16 & 96\\ 19 & 114\\ \end{array}\implies \cfrac{\frac{60}{10}+\frac{78}{13}+\frac{16}{16}}{3} \)

OpenStudy (anonymous):

@azetina